OpenStudy (anonymous):

find the points of inflection for 2x(x-4)^3

OpenStudy (anonymous):

Look at where f '' (x) = 0. f(x) = 2x (x - 4)^3 f ' (x) = 2 (x - 4)^3 + 2x * 3(x - 4)^2 f ' (x) = 2(x - 4)^3 + 6x(x - 4)^2 f '' (x) = 2*3(x - 4)^2 + 6(x - 4)^2 + 6x*2(x - 4) f '' (x) = 12(x - 4)^2 + 12x(x - 4) f '' (x) = 12(x - 4)(x - 4 + x) f '' (x) = 12(x - 4)(2x - 4) 12(x - 4)(2x - 4) = 0 (x - 4)(2x - 4) = 0 x - 4 = 0 or 2x - 4 = 0 x = 4 or 2x = 4 x = 4 or x = 2 Check x = 2: f '' (1) = 12(1 - 4)(2*1 - 4) = 12(-3)(-2) > 0 f '' (3) = 12(3 - 4)(2*3 - 4) = 12(-1)(2) < 0 Since there is a change in sign around x = 2, this is an inflection point Note that f(2) = 2*2 (2 - 4)^3 = 4(-2)^3 = 4 * -8 = -32 Check x = 4: f '' (3) = 12(3 - 4)(2*3 - 4) = 12(-1)(2) < 0 f '' (5) = 12(5 - 4)(2*5 - 4) = 12(1)(6) > 0 Since there is a change in sign around x = 4, this is an inflection point Note that f(4) = 2*4 (4 - 4)^3 = 8 * 0^3 = 8 * 0 = 0 Points of inflection: (2, -32) and (4, 0)

OpenStudy (anonymous):

Refer to a Mathematica solution attached. The results seem to be the same as @ilovenyc's .